# Unlearning Isn’t Deletion: Investigating Reversibility of Machine Unlearning in LLMs

### v1

@@ Abstract
Unlearning in large language models (LLMs) is intended to remove the influence of specific data, yet current evaluations rely heavily on token-level metrics such as accuracy and perplexity. We show that these metrics can be misleading: models often appear to forget, but their original behavior can be rapidly restored with minimal fine-tuning, revealing that unlearning may obscure information rather than erase it. To diagnose this phenomenon, we introduce a representation-level evaluation framework using PCA-based similarity and shift, centered kernel alignment, and Fisher information. Applying this toolkit across six unlearning methods, three domains (text, code, math), and two open-source LLMs, we uncover a critical distinction between reversible and irreversible forgetting. In reversible cases, models suffer token-level collapse yet retain latent features; in irreversible cases, deeper representational damage occurs. We further provide a theoretical account linking shallow weight perturbations near output layers to misleading unlearning signals, and show that reversibility is modulated by task type and hyperparameters. Our findings reveal a fundamental gap in current evaluation practices and establish a new diagnostic foundation for trustworthy unlearning in LLMs. We provide a unified toolkit for analyzing LLM representation changes under unlearning and relearning: https://github.com/XiaoyuXU1/Representational_Analysis_Tools.git.
@@ 1 Introduction
Large language models (LLMs), trained on massive corpora, have achieved remarkable success across diverse tasks, yet their capacity to memorize training snippets poses acute ethical, legal, and security risks. Memorization can unintentionally disclose sensitive, harmful, or copyrighted text [25; 13; 30], conflicting with emerging regulations, such as the EU’s Right to be Forgotten [9].
@@
Machine unlearning seeks to mitigate this threat by making a model act as though specified data were never seen [2]. Numerous methods have been proposed for LLMs [33; 12; 6; 26; 18; 17; 19; 20; 31], their success usually judged by token-level metrics, such as accuracy or perplexity.
@@
However, a pivotal question remains largely unexplored: Does LLM unlearning truly erase information, or merely suppress it, poised to “resurface” at the slightest nudge?
Empirically, many unlearning methods may just appear effective: after unlearning, a model shows near-zero accuracy or high perplexity on the forget set, but a brief fine-tuning step (even on unrelated data) can quickly restore its original behaviour (see Figure 1). This exposes a huge gap between surface-level metrics and the model’s internal state [21; 23], casting doubt on compliance and safety claims. If information is recoverable via simple fine-tuning (or relearning), can we truly claim that it has been “forgotten?” What looks like memory loss may, in fact, be a shallow perturbation.
In this work, we perform the first systematic analysis of reversibility of LLM unlearning, covering both single-shot and continual settings. The continual setting allows multiple unlearning requests over time, which we believe is a more prominent and realistic setting where the deployment of such unlearning strategies are often facing a dynamic environment. We show that standard token-level metrics prove insufficient—they can collapse even when the underlying representations remain intact. To probe deeper, we introduce a diagnostic toolkit for representational analysis, featuring PCA subspace similarity and shift [36], centered kernel alignment (CKA) [15], and Fisher information [3]. Our toolkit uncovers two distinct regimes of unlearning:
1) Reversible (catastrophic) forgetting: performance collapses, but feature subspaces are largely preserved, enabling rapid recovery, and 2) Irreversible (catastrophic) forgetting: collapse coincides with substantial representational drift, making recovery difficult/impossible. Surprisingly, both yield similar results under token-level metrics, underscoring the need for deeper representational analysis.
We further show that the transition between reversible and irreversible forgetting depends not only on the volume of unlearning requests but also on hyperparameters such as learning rate. Modest weight perturbations—especially near the output layer—can lead to token-level distortions without altering feature geometry, making “forgotten” knowledge easily recoverable.
Hence, evaluating the effectiveness of LLM unlearning must go beyond superficial token-level metrics (e.g., forget accuracy declines). In safety- and privacy-critical settings, unlearning should be judged by its ability to achieve genuine erasure rather than simply representational collapse.
@@
Figure 1: (a) Token-level accuracy and CKA subspaces of reversible (top) vs. irreversible (bottom) catastrophic forgetting due to continual unlearning then relearning, (b) Our four diagnostic tools
@@
  • We present the first systematic study of reversibility in both single and continual LLM unlearning, using a feature-space toolkit, including PCA similarity, PCA shift, CKA, and Fisher information. Our analysis distinguishes reversible from irreversible (catastrophic) forgetting.
@@
  • We conduct extensive experiments with six unlearning methods (GA [33], NPO [35], and RLabel, with their variants) across three datasets (arXiv papers, GitHub code [33], and NuminaMath-1.5 [16]) on Yi-6B [34] and Qwen-2.5-7B [32]. Our results show that standard token-level metrics (e.g., accuracy, perplexity, MIA susceptibility [27]) fail to capture true forgetting behavior.
@@
  • We theoretically analyze weight perturbations to explain how widespread vs. localized parameter changes relate to (ir)reversible forgetting. Small perturbations near the logits can distort token-level metrics despite intact features, hence leading to misleading assessments.
  • Based on our findings and extra preliminary results, we propose several future directions, including using unlearning as a complementary form of data augmentation, and designing more robust unlearning algorithms that achieve genuine forgetting while avoiding representational collapse.
2 Preliminaries
@@
LLM unlearning seeks to enhance privacy, improve safety, and reduce bias [33; 12; 26; 18; 17; 19]. Most work adopts the single-unlearning paradigm: given a training corpus $\mathcal{D}$ and a designated forget set $\mathcal{D}_{f}\subseteq\mathcal{D}$, a model $\mathcal{M}$ is first trained on $\mathcal{D}$ with algorithm $\mathcal{A}$. An unlearning procedure $\mathcal{U}$ then transforms $\mathcal{M}$ into an unlearned model $\mathcal{M}_{f}$ that should behave as if it had never encountered $\mathcal{D}_{f}$.
Ideally, $\mathcal{U}$ should produce a model statistically indistinguishable from one retrained on the retain set $\mathcal{D}_{r}=\mathcal{D}\setminus\mathcal{D}_{f}$: $\mathcal{M}_{f}=\mathcal{U}\!\bigl{(}\mathcal{M},\mathcal{D}_{f}\bigr{)}\approx\mathcal{M}_{r}=\mathcal{A}\!\bigl{(}\mathcal{M},\mathcal{D}_{r}\bigr{)}$. While current methods strike a good forget-utility balance in controlled settings [1; 8] (e.g., a fixed $\mathcal{D}_{f}$ or just a single removal request), they seldom address the practical need for continual unlearning, i.e., data owners can submit removal requests sequentially over time [1]. Let the successive forget sets be $\mathcal{D}_{f}^{(1)},\mathcal{D}_{f}^{(2)},\ldots,\mathcal{D}_{f}^{(t)}$ (whose union is $\mathcal{D}_{f}$); the retain set after $t$ rounds is $\mathcal{D}_{r}^{(t)}$. The model is then updated recursively: $\mathcal{M}_{f}^{(t+1)}=\mathcal{U}(\mathcal{M}_{f}^{(t)},\mathcal{D}_{f}^{(t+1)})$, which should be similar to $\mathcal{M}_{r}=\mathcal{A}\!\bigl{(}\mathcal{M},\mathcal{D}_{r}^{(t+1)}\bigr{)}$ at any time $t$.
Retraining LLMs is prohibitively costly, so most studies rely on empirical proxies rather than formal statistically-indistinguishable guarantees [6; 24; 18; 7]. Evaluations track forget quality on the forget set and utility on the retain set, aiming to preserve both [24]. While single unlearning often yields modest declines, it is fragile: brief fine-tuning—even on benign, unrelated data—swiftly revives the “forgotten” knowledge [1; 23; 21]. The issue worsens under continual unlearning, where each round begins from an already degraded model, ultimately triggering catastrophic forgetting—a wholesale performance collapse [1; 28]. Prior work notes this risk but does not examine its root causes.
We hypothesize that collapse does not imply true erasure; the knowledge may remain latent in the feature space. This insight leads us to distinguish two regimes of (catastrophic) forgetting:
Definition 1 (Reversible (Catastrophic) Forgetting).
Let $\theta_{0}$ denote the initial model parameters, $\mathcal{D}_{f}$ the forget set, and $\mathcal{T}$ an evaluation task with metric $E(\cdot,\mathcal{T})$. Unlearning $\mathcal{D}_{f}$ transforms the model to $\theta_{u}$. If subsequent relearning on $\mathcal{D}_{f}$ (or an equivalent reconstruction set) produces parameters $\theta_{r}$ s.t.
$E(\theta_{u},\mathcal{T})\;\ll\;E(\theta_{r},\mathcal{T})\;\approx\;E(\theta_{0},\mathcal{T}),$
the temporary performance collapse is fully reversible; we call it reversible catastrophic forgetting. When the initial degradation is modest (e.g., single unlearning), we simply call it reversible forgetting.
Definition 2 (Irreversible (Catastrophic) Forgetting).
Using the notations of Definition 1, if
$E(\theta_{u},\mathcal{T})\;\approx\;E(\theta_{r},\mathcal{T})\;\ll\;E(\theta_{0},\mathcal{T}),$
we observe irreversible catastrophic forgetting: the collapse (i.e., weight perturbation) is irreversible. We refine this to irreversible forgetting, which further requires that the irreversible degradation be restricted to the forget set; performance on the retain set and on unrelated data must remain near their original levels. This condition distinguishes targeted erasure from global model failure or collapse.
To distinguish our setting from a full retrain model, where reversibility is trivially achievable, we introduce the following restriction on the relearning phase.
Relearning Restriction. After unlearning, we briefly fine-tune $\theta_{u}$ on a small relearning set—either the cumulative forget set $\mathcal{D}_{f}=\bigcup_{t}\mathcal{D}_{f}^{(t)}$, its a distribution-similar retain set $\mathcal{D}_{r}^{(t)}$, or an unrelated out-of-distribution corpus—to obtain $\theta_{r}$ without ever revisiting the full pre-training data.
3 Token-Level Evaluation
@@ 3.1 Experiment setup
Models and Datasets. We conduct experiments on two open-source models, Yi-6B [34] and Qwen-2.5-7B [32]. To ensure the generality of our findings, we use two dataset types: i) simple tasks—arXiv abstracts and GitHub code from [33], and ii) a complex task—NuminaMath-1.5, a recent benchmark for mathematical reasoning [16]. All experiments were run on NVIDIA H100 GPUs.
Unlearning algorithms We compare six canonical methods grouped into three families.
Unlearning Scenario We consider two standard settings: i) Single unlearning: A trained model $\mathcal{M}$ receives exactly one request to remove $\mathcal{D}_{f}\subset\mathcal{D}$, producing $\mathcal{M}_{f}=\mathcal{U}(\mathcal{M},\mathcal{D}_{f})$, and ii) Continual unlearning: The model processes a stream of requests ${\mathcal{D}_{f}^{(1)},\dots,\mathcal{D}_{f}^{(k)}}$, updated iteratively by $\mathcal{M}^{(i+1)}=\mathcal{U}(\mathcal{M}^{(t)},\mathcal{D}_{f}^{(i+1)})$ with $\mathcal{M}^{(0)}=\mathcal{M}$ and $\bigcup_{i=1}^{n}\mathcal{D}_{f}^{(t)}=\mathcal{D}_{f}$. This mirrors real-world, incremental removal demands while maintaining parity with the single-step budget.
For the simple tasks, we benchmark all six algorithms: GA, GA + GD, GA + KL, RLabel, NPO, and NPO+KL. For the complex one, lacking a clearly defined retain set, we use GA, NPO, and RLabel.
Evaluation Metrics For single unlearning (simple tasks only), we report: forget-set accuracy (F.Acc), retain-set accuracy (R.Acc), and privacy leakage via min-$k\%$-prob MIA AUC [27].
@@
Relearning setting. To gauge how readily forgotten knowledge re-emerges, each unlearning run is followed by a controlled relearning phase. Single unlearning: We fine-tune on the whole forget set $\mathcal{D}_{f}$ once, producing a single-step relearned model. Continual unlearning: For settings that trigger catastrophic collapse, we fine-tune with three cases: i) the cumulative forget set $\bigcup_{i}\mathcal{D}_{f}^{(t)}$, ii) the corresponding retain set $\mathcal{D}_{r}^{(t)}$, and iii) an unrelated auxiliary corpus. These progressively relax assumptions about access to the forgotten content, revealing the recovery potential in each case.
Hyperparameter Configuration. To comprehensively evaluate the effects of unlearning, we design multiple hyperparameter configurations that vary both the learning rate and the number of unlearning requests. For single unlearning we sweep the learning rate over $LR\in\{3,4,5\}\!\times\!10^{-6}$ while fixing the request count to $N=1$. For continual unlearning we vary both knobs: on the simple task (Yi-6B) we test $LR\in\{3,5\}\!\times\!10^{-6}\cup\{3\times 10^{-5}\}$ with $N\in\{6\to 100\}$; on the complex task (Qwen-2.5-7B) we use $LR\in\{3,5\}\!\times\!10^{-6}$ and $3\times 10^{-5}$ together with $N\in\{6\to 100\}$. All runs adopt the optimizer settings of [1]: AdamW [22] ($\beta_{1}=0.9$, $\beta_{2}=0.95$, $\varepsilon=10^{-8}$), a cosine schedule with 10% warm-up followed by decay to 10% of peak, weight decay 0.1, and gradient clipping at 1.0.
Phase Method LR=3$\times$10-6 LR=4$\times$10-6 LR=5$\times$10-6
@@
3.2 Token-Level Evaluation Results
We report quantitative results on both single unlearning and continual unlearning settings using Yi-6B and Qwen-2.5-7B across multiple configurations (Tables 13). For completeness, detailed experimental results and additional settings are provided in Appendix A.4.
Single Unlearning. For Yi-6B under single unlearning, Table 1 shows that all six unlearning methods reduce MIA and F.Acc, indicating some level of unlearning. The changes on retain set are modest: R.Acc only drops 2–5% for most methods, and MIA decreases by less than 30 points in most cases. Importantly, the relearned models often recover original performance—both GA+KL and RLabel restore R.Acc to near 65.0%, and F.Acc rebounds above 77%. These results suggest that under single unlearning, most methods achieve seemingly successful forgetting at the token level, but as we show in Section 4.2 to Section 4.2, the underlying representation changes are minimal—indicating the phenomenon of reversible forgetting.
Continual Unlearning. By examining post-relearning recoverability in Table 2 and Table 3, we identify two distinct forms of catastrophic forgetting. When the model regains both utility (e.g., F.Acc, R.Acc) and privacy (e.g., MIA AUC) to levels near or exceeding the original after relearning, we classify the behavior as reversible catastrophic forgetting. This suggests that the underlying representational structure remains intact, enabling efficient recovery via lightweight retraining. Such reversibility is consistently observed in methods like NPO and NPO+KL, particularly under low learning rates or small removal batches.
Conversely, when relearning fails to restore utility—reflected in persistently low F.Acc and R.Acc despite partial MIA recovery—we categorize it as irreversible catastrophic forgetting. This scenario frequently arises with methods like GA and RLabel under aggressive hyperparameters (e.g., $\mathrm{LR}=3\times 10^{-5}$), where damage accumulates across layers and results in irreversible representational collapse. Importantly, MIA AUC alone can be misleading. Models may exhibit near-complete privacy recovery while remaining functionally impaired.
Table 2: Yi-6B: MIA / F.Acc / R.Acc (%) for simple task under four unlearning settings
Phase Method LR=3$\times$$10^{-5}$, N=100 LR=5$\times$$10^{-6}$, N=100 LR=3$\times$$10^{-6}$, N=100 LR=3$\times$$10^{-5}$, N=6

### v2

@@ Abstract
Unlearning in large language models (LLMs) aims to remove specified data, but its efficacy is typically assessed with task-level metrics like accuracy and perplexity. We demonstrate that these metrics are often misleading, as models can appear to forget while their original behavior is easily restored through minimal fine-tuning. This phenomenon of reversibility suggests that information is merely suppressed, not genuinely erased. To address this critical evaluation gap, we introduce a representation-level analysis framework. Our toolkit comprises PCA-based similarity and shift, centered kernel alignment (CKA), and Fisher information, complemented by a summary metric, the mean PCA distance, to measure representational drift. Applying this framework across six unlearning methods, three data domains, and two LLMs, we identify four distinct forgetting regimes based on their reversibility and catastrophicity. Our analysis reveals that achieving the ideal state–irreversible, non-catastrophic forgetting–is exceptionally challenging. By probing the limits of unlearning, we identify a case of seemingly irreversible, targeted forgetting, offering new insights for designing more robust erasure algorithms. Our findings expose a fundamental gap in current evaluation practices and establish a representation-level foundation for trustworthy unlearning. [^1] https://github.com/XiaoyuXU1/Representational_Analysis_Tools
@@ 1 Introduction
Large language models (LLMs), trained on massive corpora, have achieved remarkable success across diverse tasks, yet their capacity to memorize training snippets poses acute ethical, legal, and security risks. Memorization can unintentionally disclose sensitive, harmful, or copyrighted text [28; 14; 35], conflicting with emerging regulations, such as the EU’s Right to be Forgotten [8].
@@
Machine unlearning seeks to address this challenge by algorithmically erasing the influence of specified data, making a model behave as if it had never been trained on that data [2]. While numerous unlearning methods have been developed for LLMs [40; 12; 29; 20; 19; 22; 37], their efficacy is typically assessed using task-level metrics, such as accuracy on a held-out “forget set.”
@@
However, these evaluations overlook a pivotal question: Does LLM unlearning achieve genuine erasure, or does it merely suppress information, leaving it poised to resurface?
Emerging evidence suggests that many unlearning methods may only be superficially effective. For instance, after unlearning, a model may exhibit poor performance on the forget set, yet this “forgotten” knowledge can often be rapidly restored through a brief fine-tuning step, even on unrelated data [24; 26] (see Figure 1). Similarly, low-bit quantization of model weights has been shown to recover forget-set performance without any retraining, offering a simple probe for this reversibility [44]. These findings reveal a critical gap between task-level metrics and the model’s internal state. If purportedly erased information can be easily recovered, the unlearning process may amount to little more than a shallow perturbation, failing to provide meaningful safety or compliance guarantees.
While prior work has noted the phenomenon of reversibility [44], it has not been systematically investigated, particularly in the context of continual unlearning, a more realistic scenario where deletion requests arrive sequentially over a model’s lifecycle. This paper presents the first systematic analysis of the reversibility of LLM unlearning. We argue that standard task-level metrics are insufficient, as performance can collapse (a.k.a. catastrophic forgetting) even when the underlying feature representations remain largely intact. To move beyond superficial evaluations, we introduce a diagnostic toolkit for representational analysis, which includes PCA subspace similarity and shift [45], centered kernel alignment (CKA) [16], and Fisher information (FIM) [3]. Building on this, we propose the mean PCA distance as a quantitative measure to summarize the degree of representational drift.
With our toolkit, we build a taxonomy characterizing unlearning along two axes: reversibility and catastrophicity (collateral damage to retained knowledge). This allows us to distinguish four regimes: 1) Reversible, Catastrophic: Global performance collapse that is fully recoverable via relearning.
2) Reversible, Non-Catastrophic: Targeted performance modest degradation that is easily restored.
3) Irreversible, Catastrophic: Permanent and unrecoverable global performance collapse.
4) Irreversible, Non-Catastrophic: Ideally, permanent erasure of target data without collateral damage.
Crucially, we find that task-level metrics (e.g., forget accuracy) alone cannot distinguish between reversible and irreversible (catastrophic) forgetting, underscoring the necessity of our representation-level analysis; the evaluation must transcend task-level performance. We further show that transitions between these regimes depend on the unlearning method, the volume of unlearning requests, and key hyperparameters. In safety-critical applications, the central criterion for success should be the demonstrable achievement of irreversible erasure, not merely superficial performance degradation.
@@
Figure 1: (a) task-level accuracy and CKA subspaces of reversible (top) vs. irreversible (bottom) catastrophic forgetting due to continual unlearning then relearning, (b) Our four diagnostic tools
@@
  • We present the first systematic study of reversibility in both single and continual LLM unlearning. We introduce a representation-level diagnostic toolkit and a quantitative metric, the mean PCA distance, to analyze representational drift and distinguish four regimes of forgetting.
@@
  • We conduct extensive experiments with six unlearning methods on two LLMs (Yi-6B and Qwen-2.5-7B) across three distinct domains. Our results demonstrate that standard task-level metrics (e.g., accuracy, perplexity, MIA susceptibility) are insufficient for assessing the true extent of unlearning.
@@
  • We theoretically analyze weight perturbations to explain how widespread vs. localized parameter changes relate to (ir)reversible forgetting. Small perturbations near the logits can distort task-level metrics despite intact features, hence leading to misleading assessments.
  • We identify a case of seemingly irreversible, non-catastrophic forgetting, offering insights for designing more robust unlearning algorithms. We also highlight the potential for unlearning to serve as a form of data augmentation, improving model representations upon relearning.
2 Preliminaries and Our Formulation
@@
LLM unlearning aims to remove the influence of specific data from a trained model to enhance privacy, improve safety, or mitigate bias [40; 12; 29; 20; 19; 22]. The standard paradigm involves a training corpus $\mathcal{D}$, from which $\mathcal{D}_{f}\subseteq\mathcal{D}$ is designated as the forget set. A model $\mathcal{M}$ is first trained on $\mathcal{D}$ via an algorithm $\mathcal{A}$. An unlearning procedure $\mathcal{U}$ then transforms $\mathcal{M}$ into unlearned $\mathcal{M}_{f}$, which should ideally behave as if it were trained only on the retain set $\mathcal{D}_{r}=\mathcal{D}\setminus\mathcal{D}_{f}$. Formally, the goal is to statistically approximate the retrained model $\mathcal{M}_{r}$: $\mathcal{M}_{f}=\mathcal{U}(\mathcal{M},\mathcal{D}_{f})\approx\mathcal{M}_{r}=\mathcal{A}(\mathcal{D}_{r}).$
Retraining LLMs is prohibitively costly, so most studies rely on empirical proxies rather than formal statistically-indistinguishable guarantees [27; 20; 6]. Evaluations track forget quality on the forget set, utility, and downstream task accuracy on the retain set, aiming to preserve both dimensions.
While current methods can achieve reasonable balances between forgetting and utility in single-shot scenarios [1; 7], they often falter in the practical continual setting, where removal requests arrive sequentially over time [1]. For a sequence of forget sets $\mathcal{D}_{f}^{(1)},\mathcal{D}_{f}^{(2)},\ldots,\mathcal{D}_{f}^{(t)}$ (the union is $\mathcal{D}_{f}$), the retain set is $\mathcal{D}_{r}^{(t)}$ after $t$ rounds. The model is then updated recursively: $\mathcal{M}_{f}^{(t)}=\mathcal{U}(\mathcal{M}_{f}^{(t-1)},\mathcal{D}_{f}^{(t)})$, which should be similar to $\mathcal{M}_{r}=\mathcal{A}\!\bigl(\mathcal{M},\mathcal{D}_{r}^{(t)}\bigr),\forall t$. However, empirically, it often leads to catastrophic forgetting–a severe decline in performance on both forgotten and retained knowledge [1; 32].
Single-shot unlearning is “fragile:” fine-tuning, even on benign, unrelated data, can rapidly restore the supposedly “forgotten” knowledge [1; 26; 24]. Such fragility persists in continual unlearning as well. Prior work has noted this phenomenon but has not deeply investigated its underlying mechanics.
2.1 A Taxonomy of Forgetting Regimes
We hypothesize that this performance collapse does not necessarily equate to true information erasure; the knowledge might merely become latent or suppressed. To formalize this hypothesis, we introduce a taxonomy of forgetting based on two axes: catastrophicity (the extent of collateral damage to retained knowledge) and reversibility (whether forgotten knowledge can be recovered).
Let $\theta_{0}$ be the initial model parameters, $\theta_{u}$ be the parameters after unlearning, and $\theta_{r}$ be the parameters after a subsequent relearning phase (defined below). We use $E(\theta,\mathcal{T})$ to denote a performance metric (e.g., accuracy) evaluated on a task set $\mathcal{T}$, which can be partitioned into a forget-related task $\mathcal{T}_{f}$ and a retain-related task $\mathcal{T}_{r}$. We define four distinct regimes of forgetting, summarized in Table 1.
Definition 1 (Four Regimes of Forgetting).
Let $\Delta_{u}(\mathcal{T})=E(\theta_{0},\mathcal{T})-E(\theta_{u},\mathcal{T})$ be the performance drop after unlearning, and $\Delta_{r}(\mathcal{T})=E(\theta_{0},\mathcal{T})-E(\theta_{r},\mathcal{T})$ be the residual drop after relearning. The nature of forgetting is determined by these drops on the forget set ($\mathcal{T}_{f}$) and retain set ($\mathcal{T}_{r}$).
Catastrophic vs. Non-Catastrophic: Forgetting is catastrophic if utility on the retain set degrades significantly (both $\Delta_{u}(\mathcal{T}_{r})\ and\ \Delta_{u}(\mathcal{T}_{u})\gg 0$) and non-catastrophic otherwise ($\Delta_{u}(\mathcal{T}_{r})\approx 0$).
Reversible vs. Irreversible: Forgetting is reversible if relearning almost recovers initial performance on forget set ($\Delta_{r}(\mathcal{T}_{f})\approx 0$) and irreversible if a significant performance drop persists ($\Delta_{r}(\mathcal{T}_{f})\gg 0$).
The combination of these two properties yields four regimes, among which the irreversible, non-catastrophic forgetting is deemed ideal, but remains challenging to achieve in practice.
Relearning Restriction. To test reversibility, we must define a constrained relearning process that is distinct from a full retrain. After unlearning yields $\theta_{u}$, we obtain $\theta_{r}$ by briefly fine-tuning $\theta_{u}$ on a small relearning set, without access to raw pre-training data. This set is size-matched to $\mathcal{D}_{f}$ (i.e., total budget is $|\mathcal{D}_{f}|$) and drawn from one of three sources: (i) the forget set $\mathcal{D}_{f}$ itself, representing a worst-case recovery scenario, (ii) a retain subset $\mathcal{D}_{r}^{(t)}$ from domains similar to $\mathcal{D}_{f}$, or (iii) general-domain, out-of-distribution data. This ensures a fair and direct assessment of knowledge recoverability.
[…]
Regime Observed Description
[…]
3 Classic (Task-Level) Evaluation Can Be Deceptive
@@ 3.1 Experiment setup
Models and Datasets. We adopt two open-source LLMs: Yi-6B [41] and Qwen-2.5-7B [39]. To assess the generality of our findings, we employ two distinct dataset types for unlearning: (i) simple tasks, comprising arXiv abstracts and GitHub code from [40], and (ii) a complex task, NuminaMath-1.5, a recent benchmark for mathematical reasoning [18]. All experiments are performed on NVIDIA H100 GPUs. (Additional results on TOFU [27] and the Traditional-Chinese corpus [^2] https://huggingface.co/datasets/taide/taide-bench are in Appendix.)
@@
Unlearning algorithms. We evaluate six canonical unlearning methods, organized into three families.
Unlearning Scenarios. We consider two standard settings: i) Single unlearning: A trained model $\mathcal{M}$ receives exactly one request to remove $\mathcal{D}_{f}\subset\mathcal{D}$, and ii) Continual unlearning: The model processes a stream of requests $\mathcal{D}_{f}^{(1)},\dots,\mathcal{D}_{f}^{(t)}$}, yielding a sequence of models where $\mathcal{M}^{(t)}=\mathcal{U}(\mathcal{M}^{(t-1)},\mathcal{D}_{f}^{(t)})$.
For simple tasks, we benchmark all six algorithms. For the complex math reasoning task, where a well-defined retain set is not available, we evaluate the core GA, NPO, and RLabel methods.
Evaluation Metrics. In single unlearning (simple tasks), we measure forget-set accuracy (F.Acc), retain-set accuracy (R.Acc), and privacy leakage via min-$k\%$-prob MIA AUC [31].
[…]
Relearning Setting. To assess the reversibility and of unlearning, each run is followed by a controlled relearning phase. The unlearned model is briefly fine-tuned on specific data without access to the full pre-training corpus. For single unlearning, we fine-tune once on the entire forget set $\mathcal{D}_{f}$. For continual unlearning, we evaluate three conditions: (i) the cumulative forget set $\bigcup_{t}\mathcal{D}_{f}^{(t)}$, representing a worst-case adversarial scenario, (ii) the corresponding retain subset $\mathcal{D}_{r}^{(t)}$, as a proxy for the data distribution, and (iii) unrelated out-of-distribution data (general-domain samples explicitly different from $\mathcal{D}_{f}$). Each relearning dataset is size-matched to its corresponding unlearning request.
Hyperparameter Configuration. To comprehensively evaluate the effects of unlearning, we design multiple hyperparameter configurations that vary both the learning rate and the number of unlearning requests. For single unlearning we sweep the learning rate over $\mathrm{LR}\in\{3,4,5\}\!\times\!10^{-6}$ while fixing the request count to $N=1$. For continual unlearning we vary both knobs: on the simple task (Yi‑6B) we test $\mathrm{LR}\in\{3,5\}\!\times\!10^{-6}\cup\{3\times 10^{-5}\}$ with $N\in\{6\to 100\}$; on the complex task (Qwen‑2.5‑7B) we use $\mathrm{LR}\in\{3,5\}\!\times\!10^{-6}$ and $3\times 10^{-5}$ together with $N\in\{6\to 100\}$. All runs adopt the optimizer settings of [34]: AdamW [25] ($\beta_{1}=0.9$, $\beta_{2}=0.95$, $\varepsilon=10^{-8}$), a cosine schedule with 10% warm‑up followed by decay to 10% of peak, weight decay 0.1, and gradient clipping at 1.0.
3.2 Evaluation Results
We report quantitative results for single and continual unlearning on Yi-6B and Qwen-2.5-7B under various configurations. Complete results are provided in Appendix Tables 6 and 7.
Single Unlearning. On Yi-6B, all six methods successfully reduce MIA and F.Acc, indicating a certain degree of forgetting (Table 2). The impact on the retain set is modest, with R.Acc dropping by only 2–5%. However, relearning often restores original performance; for instance, GA+KL and RLabel recover R.Acc to approximately 65% and F.Acc above 77%. These findings suggest that single unlearning achieves superficial forgetting, as the underlying representations remain largely intact (Section 4.2.1). This outcome characterizes the reversible, non-catastrophic forgetting regime.
Table 3: Yi-6B: MIA / F.Acc / R.Acc (%) for simple task under four unlearning settings. Bold numbers indicate improvements over the Original baseline in F.Acc or R.Acc.
[…]